OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with slope -2/3 that passes through (-2, 5).

OpenStudy (anonymous):

First, use the point-slope form for a line:\[y - y _{1} = m(x - x _{1})\] where "m" is your given slope, and (x1, y1) is either of the given points. Then, rearrange into form: y= mx + b